OpenStudy (anonymous):

What is the Celsius temperature of 1 mole of a gas that has an average kinetic energy of 4,265 joules? 69°C 96°C 369°C 435°C @abhisar

OpenStudy (anonymous):

@RosieF

OpenStudy (anonymous):

@Abmon98

OpenStudy (abhisar):

\(\tt \huge E_k = \frac{3}{2} R T\)

OpenStudy (abhisar):

T should be in kelvin, R is boltzman constant

OpenStudy (abhisar):

Can u calculate it now ?

OpenStudy (anonymous):

no am slow in chem @Abhisar

OpenStudy (abhisar):

do you know the value of boltzman constant ?

OpenStudy (abmon98):

@veronica.montanez i have no idea i am sorry.

OpenStudy (anonymous):

so u may not know sometging simliary would you @Abmon98 and no @Abhisar

OpenStudy (abhisar):

it is 8.3145 J/(mol K)

OpenStudy (abhisar):

now just plug in the values into the equation i just gave to u and find the value of T

OpenStudy (abmon98):

no

OpenStudy (anonymous):

so I would be 4265/8.3145*3/2

OpenStudy (abhisar):

yes , find the value

OpenStudy (anonymous):

how ? it gave me 769 its not the answer @Abhisar

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @Abhisar \(\tt \huge (4265\times2)/3= R T\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

---> 2843/8.3145 = T = 341.93 K

OpenStudy (abhisar):

T = 68.93 = 69°

OpenStudy (anonymous):

oo ok could u help me on two more @Abhisar

OpenStudy (abhisar):

I would love to do that !

OpenStudy (abhisar):

but it's 2:23 AM here !

OpenStudy (anonymous):

oo sorry @Abhisar

OpenStudy (abhisar):

It's ok ! u can post your questions....if someone other helps you then it's cool or else i'll check it some times after :)

OpenStudy (anonymous):

ok goomorning/night @Abhisar